# Gradient Descent on Logistic Regression with Non-Separable Data and Large Step Sizes

### v1

@@ 5 One-dimensional case
resulting in a cycle. As shown in the second panel of Figure 4, we have effectively added enough gradient to the left of the minimizer such that $\tilde{w}_{k+1}$ gets kicked right back to where we started.
However, adding a ReLU to a loss function of the form 7 does not automatically yield a valid classification problem. We need to show that this can be achieved equivalently by adding more examples to the dataset. Recall that by Corollary 1, this ReLU is the limit of the function $\ell_{\epsilon}(w)=\epsilon^{2}\ell\left(-w/\epsilon^{2}\right)$ as $\epsilon\to 0$ (Figure 3 and Corollary 1). This allows us to define a continuous perturbation
${\mathcal{L}}_{\epsilon}(w)={\mathcal{L}}(w)+\frac{w_{0}-w_{k+1}}{\eta}\ell_{\epsilon}(w).$ (12)
It remains to invoke the implicit function theorem to argue that there exists a non-zero $\epsilon$ such that $\epsilon^{2}(w_{0}-w_{k+1})/\eta>0$ is rational. This implies we can obtain a loss function equivalent to adding integer copies of the example $x=1/\epsilon^{2}$ of label $y=1$ to the original dataset. This ensures that ${\mathcal{L}}_{\epsilon}$ indeed corresponds to an objective on a valid binary classification dataset.
Finally, we argue that this cycle is locally stable. This can be achieved by observing that at all points on this cycle (except for the one to the left of $w^{*}$) have a Hessian strictly smaller than that at $w^{*}$ (Assumption 1). This is sufficient to guarantee that the Jacobian of the $k+1$-th iterate GD map has a magnitude strictly less than one (if this is not the case, we can extend the trajectory backwards from $w_{0}$ to $w_{-1}$, $w_{-2}$, …, until satisfied). Local stability implies that if we initialize sufficiently close to the cycle (rather than the minimizer), GD would converge to this cycle instead.
@@
As verifying the assumptions of the implicit function theorem can be technical and tedious, we defer the full proof to Section A.2. In Figure 5, we illustrate that one can easily construct a dataset for different values of $\gamma\in(1,2]$ such that GD converges to cycles. A pattern to note is that the smaller the $\gamma$, the longer period the cycle tends to have, as it takes more steps to move along the “flatter” side of the loss before being kicked back. Note that convergence to $w^{*}$ and stable cycles are not necessarily the only two events that can happen below the $2/\lambda$ step size. If for some dataset there exists a period-$3$ cycle under the step size $\eta=\gamma/\lambda$ for some $\gamma<2$, then a chaotic GD trajectory can also occur for that dataset under the same $\gamma$ (Li and Yorke, 1975).
Figure 5: For each $\gamma$, we construct a one-dimensional dataset such that GD on this problem with the logistic loss converges to a stable cycle under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$. One exception is the last column which seems to suggest GD can even be chaotic when $\gamma<2$. Figures in the first row show the loss evaluated at successive iterates, while the second row shows the power spectral density of the losses at the last $1024$ iterations. Details on each dataset are in Appendix E.
@@ 6 Higher dimensions
In the one-dimensional case, we have shown that under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$, convergence is guaranteed for all $\gamma\leq 1$, while increasing $\gamma$ beyond $1$ can lead to cycles. In higher dimensions, we no longer have global convergence even in the small $\gamma$ case. As we now show, for all $\gamma\in(0,1]$, a cycle can be constructed in $d=2$ using similar techniques. Since we can embed the constructed dataset in higher dimensions, the result trivially extends to any non-linearly-separable classification problem satisfying Assumption 1 of dimension $d>1$.
@@
Suppose Assumption 1 holds and $d=2$. Then for all $\gamma\in(0,1]$, there exists a non-separable classification problem on which a GD trajectory under the step size $\eta=\gamma/\lambda$ converges to a cycle of period $k>1$, where $\lambda=\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$.
@@
Fix a $\gamma\in(0,1]$, we construct a base dataset corresponding to a loss ${\mathcal{L}}$ such that the minimizer is strictly positive. We show that there exists a trajectory that moves almost in a straight line towards the minimizer, as in the first plot of Figure 6. Let this trajectory be $w_{0},w_{1},\dots$, up to $w_{k}$, and $T(w)=w-\eta\nabla{\mathcal{L}}(w)$ be the GD update on ${\mathcal{L}}$.
@@
for some constants $c_{1},c_{2}>0$ and vectors $p$ and $q$. Specifically, $p$ and $q$ are positioned such that the minimizer of ${\mathcal{L}}_{\text{perturbed}}$ is identical to that of ${\mathcal{L}}$. Let $\left\{\tilde{w}_{t}\right\}$ be the trajectory of running GD on ${\mathcal{L}}_{\text{perturbed}}$ with the step size $\eta=\gamma/\lambda_{\max}(\nabla^{2}{\mathcal{L}}_{\text{perturbed}}(w^{*}))=\eta$, and the same initialization $\tilde{w}_{0}=w_{0}$.
Observe that the two ReLUs are activated for all points above the line $w^{\mathsf{\scriptscriptstyle T}}p=0$ and below the line $w^{\mathsf{\scriptscriptstyle T}}q=0$. We design our original trajectory such that neither gets activated until $w_{k-1}$, thus the two trajectories coincide for all $w_{t}$ from $t=0$ to $t=k-2.$ At $\tilde{w}_{k-1}=w_{k-1}$, we cross over $w^{\mathsf{\scriptscriptstyle T}}p=0$, the first ReLU is activated, kicking the perturbed trajectory in the orthogonal direction to $\tilde{w}_{k}$ (as in the middle plot of Figure 6). At this new point $\tilde{w}_{k}$, we carefully set $c_{1}$, $c_{2}$, $p$ and $q$ such that the first ReLU becomes inactive while the second activates, and that another GD step takes us back to $w_{0}$.
As in the one-dimensional case (Theorem 2), we then use the fact that ${\mathcal{L}}_{\text{perturbed}}$ is the limit of a continuous perturbation for which we can find a valid classification dataset. Stability of the resulting cycle is also straightforward.
Figure 6: Illustration of our cycle construction in two dimensions using the logistic loss. The objective (in log scale) in the middle and right contours are created by adding large-magnitude examples to the dataset underlying ${\mathcal{L}}$, depicted on the left. Red stars mark the minimizer $w^{*}$ of the respective losses. Starting at the same $w_{0}$ using the same $\gamma=0.4$, GD on ${\mathcal{L}}$ and ${\mathcal{L}}$ with one set of additional data points both converge to the solution. When we add in the second set of data points, GD converges to a period-$13$ cycle. The trajectory in white marks the first $500$ steps, while the final $20$ steps are marked in red. Details on the synthetic dataset are in Appendix E.
@@
Suppose Assumption 1 holds and $d>1$. Then for all $\gamma\in(0,1]$, there exists a non-separable classification problem on which a GD trajectory under the step size $\eta=\gamma/\lambda$ converges to a cycle of period $k>1$, where $\lambda=\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$.
@@
By Theorem 3, there exists a dataset in two dimensions such that the statement holds. Let $Z\in\mathbb{R}^{n\times 2}$ and label vector $y$ be this dataset, and $u^{*}\in\mathbb{R}^{2}$ be the minimizer of the corresponding loss. We lift this dataset to higher dimensions by simply padding 00’s to each example $x_{i}$ in $X$ until it is a vector of dimension $d$. Let $x_{i}=(z_{i,1},z_{i,2},0,\dots,0)\in\mathbb{R}^{d}$ where $d>2$ for each $i=1,\dots,n$, and $w^{*}=(u^{*}_{1},u^{*}_{2},0,\dots,0)\in\mathbb{R}^{d}$. Observe that $w^{*}$ is a minimizer of the loss on the new dataset $(X,y)$, because the loss is convex and
$\nabla{\mathcal{L}}(w^{*})=\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(-y_{i}x_{i}^{\mathsf{\scriptscriptstyle T}}w^{*})(-y_{i}x_{i})=\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(-y_{i}z_{i}^{\mathsf{\scriptscriptstyle T}}u^{*})(-y_{i}x_{i})$
which is just $\nabla{\mathcal{L}}(u^{*})=0$ but padded with more zeros. Similarly, we also have $\lambda_{\max}(\nabla^{2}{\mathcal{L}}(u^{*}))=\lambda_{\max}(\nabla^{2}{\mathcal{L}}(w^{*}))$, thus the step sizes used on the two datasets are the same under the same $\gamma$. Therefore, we can take the cycle in the two-dimensional dataset, pad each point on the orbit by 00’s to obtain a cycle in the higher dimension. ∎
@@ Appendix A Proofs in one dimension
For all proofs in this section, we assume without loss of generality that $y_{i}=1$ for all $i$. If $y_{i}=-1$ for some $i$, we can simply flip the corresponding $x_{i}$ to $-x_{i}$ to obtain the same objective. We use $T$ to denote the GD map with a constant step size $\eta$
@@ A.1 Convergence under the stable step size
We assume without loss of generality that $w^{*}>0$. If $w^{*}=0$, then fact that ${\mathcal{L}}^{\prime\prime}$ is maximized at 00 implies $T$ is Lipschitz with a constant strictly less than $1$ (because ′′${\mathcal{L}}^{\prime\prime}(w^{*})>0$ when the data is non-separable). Therefore $T$ is a contraction which implies global convergence. If $w^{*}<0$, we can flip the signs of all the $x_{i}$’s and get $w^{*}>0$ as the problem is symmetric about 00. The proof is split into three cases based on the initialization:
@@
$=\underbrace{w^{*}-\eta C}_{\bar{w}}$ (17)
for some constant $C$ due to boundedness of $\ell^{\prime}$. As $\bar{w}$ does not depend on $t$, we can never move arbitrarily far to the right. It remains to establish the iteration complexity.
Case 2: $w_{0}<0$. In this case, the maximum number of iterations for us to arrive at $w_{\tau}\geq 0$ is given by
@@
which holds for all $\eta>0$. This is because while ${\mathcal{L}}^{\prime}<0$, it is decreasing in magnitude as we increase from $w_{0}$ to 00.
@@
So in this interval, we cross over into ${\mathcal{I}}$ in one step. Combining 16, 18 and A.1, we have that after at most $1+\max\left\{0,\,-w_{0}\cdot{\mathcal{L}}^{\prime\prime}(w^{*})/{\mathcal{L}}^{\prime\prime}(0)\right\}$ number of iterations, we will enter ${\mathcal{I}}$, from which convergence is guaranteed at a linear rate. The rate depends on $\eta=1/{\mathcal{L}}^{\prime\prime}(w^{*})$ as well as $\bar{w}$ if we did not initialize in ${\mathcal{I}}$. If $w^{*}<0$, then the initial number of iterations kick in when $w_{0}>0$ and the entire argument holds with a sign flip. This completes the proof. ∎
@@ A.2 Cycle construction below the critical step size
$w_{k}<0<w_{k-1}<\dots<w_{1}<w_{0}$ (19)
is possible for some $k>2$, under the step size $\eta$ where $w_{t+1}=T(w_{t})$ for $t=0,1,\dots,k-1$. We also assume that $w_{0}>T(w_{k})$. This is valid since if it’s not the case, we can use Lemma 2 to extend the trajectory further to the right as $w_{0}<w_{-1}<\dots<w_{-t}$ until $w_{-t}>T(w_{k})$ is satisfied, then relabel the iterations.
@@
$\ell_{\epsilon}(w)\coloneqq\begin{cases}\epsilon^{2}\cdot\ell\left(-\frac{w}{\epsilon^{2}}\right)&\text{if $\epsilon\neq 0$}\\ \max\{-w,\,0\}&\text{if $\epsilon=0$}\end{cases}$ (20)
where $a$ is the upper bound on $\ell^{\prime}$ given in Assumption 1. Define constant
$c\coloneqq\frac{w_{0}-T(w_{k})}{\eta}>0,$ (21)
@@
be the corresponding GD map. Starting at $\tilde{w}_{0}=w_{0}$ in 19, let $\tilde{w}_{t+1}=T_{\epsilon}(\tilde{w}_{t})$ for $t=0,1\dots,k$. Since $w^{*}_{\epsilon}$, ${\mathcal{L}}^{\prime}_{\epsilon}$, and ${\mathcal{L}}^{\prime\prime}_{\epsilon}$ are all continuous in $\epsilon$ (Lemmas 3 and 4), so is $\eta_{\epsilon}$ and $T_{\epsilon}$, with limits
@@
as $\epsilon\to 0$. Furthermore, for all $w>0$, $\lim_{\epsilon\to 0}\ell^{\prime}_{\epsilon}(w)=0$. Combined with the fact that original trajectory 19 is positive up to $w_{k-1}$, we also have $\tilde{w}_{t}\to w_{t}$ as $\epsilon\to 0$ for all $t=0,1,\dots,k$. We will show that the next step takes us back to $w_{0}$, resulting in a cycle.
@@
$G(\epsilon,w)\coloneqq T^{k+1}_{\epsilon}(w)-w=\underbrace{T_{\epsilon}\circ\ldots\circ T_{\epsilon}}_{k+1\text{ times}}(w)-w.$ (24)
@@
$G(0,w_{0})$ $=T_{0}(T^{k}_{0}(w_{0}))-w_{0}$
$=T_{0}\left(w_{k}\right)-w_{0}$
$=w_{k}-\eta L^{\prime}(w_{k})+\eta\cdot c-w_{0}$ ($w_{k}<0$)
$=T(w_{k})+w_{0}-T(w_{k})-w_{0}$ (Definition of $c$ in 21)
@@
$\frac{\partial G}{\partial w}(0,w_{0})$ $=\prod_{t=0}^{k}\left(1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(T_{0}(\tilde{w}_{t}))\right)-1.$
Note that for all $t=0,1,\dots,k-1$, $\tilde{w}_{t}=w_{t}>w^{*}$ when $\epsilon=0$ and that $w^{*}=w^{*}_{0}$,
@@
$\left|\prod_{t=0}^{k-1}\left(1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(T_{0}(\tilde{w}_{t}))\right)\right|<\left|(1-\gamma)^{k}\right|\leq 1$
@@
$\mu(\epsilon,w)$ $\coloneqq\left|(T_{\epsilon}^{k+1})^{\prime}(w)\right|=\left|\prod_{t=0}^{k}(1-\eta_{\epsilon}{\mathcal{L}}_{\epsilon}^{\prime\prime}(T^{t}_{\epsilon}(w)))\right|,$ (25)
@@
We know that $\mu(0,w_{0})<1$ from earlier, and by continuity there must exist a neighbourhood $U$ around $\epsilon=0$ such that $\mu(\epsilon,\omega(\epsilon))<1$ for all $\epsilon\in U$, as $\omega$ is continuous in $\epsilon$. Since $U$ and $I$ are both open, there must exist a nonzero $\epsilon_{0}\in I\cap U$ such that
@@
implying that the trajectory on the perturbed objective with $\epsilon=\epsilon_{0}$ is a stable cycle. Stability ensures that if initialize close enough to $\omega(\epsilon_{0})$ we will converge to this cycle, as it is a stable fixed point of the $T^{k+1}$ map. A final note is that we can always pick $\epsilon_{0}$ such that $c\cdot\epsilon_{0}^{2}$ is rational, so that the perturbed objective 22 indeed corresponds to a valid classification problem. This completes the proof. ∎
Lemma 1 (Crossing over $w^{*}$).

### v2

@@ 5 One-dimensional case
resulting in a cycle. As shown in the second panel of Figure 4, we have effectively added enough gradient to the left of 00 such that $\tilde{w}_{k+1}$ gets kicked right back to where we started.
@@
However, adding a ReLU to a loss function of the form 7 does not automatically constitute a valid classification problem. We need to show that the kick can be achieved equivalently by adding more examples to the dataset, yielding an objective of the finite-sum form Equation 7. Recall that by Corollary 1, this ReLU is the limit of the function $\ell_{\epsilon}(w)=\epsilon^{2}\ell\left(-w/\epsilon^{2}\right)$ as $\epsilon\to 0$ (Figure 3 and Corollary 1). This allows us to define a continuous perturbation
${\mathcal{L}}_{\epsilon}(w)={\mathcal{L}}(w)+\frac{w_{0}-w_{k}}{\eta}\ell_{\epsilon}(w).$ (12)
It remains to invoke the implicit function theorem to argue that there exists a non-zero $\epsilon$ such that $\epsilon^{2}(w_{0}-w_{k})/\eta>0$ is rational. This implies we can obtain a loss function equivalent to adding integer copies of the example $x=1/\epsilon^{2}$ of label $y=1$ to the original dataset. This ensures that ${\mathcal{L}}_{\epsilon}$ indeed corresponds to an objective on a valid binary classification dataset.
Finally, we argue that this cycle is locally stable. This can be achieved by observing that ${\mathcal{L}}^{\prime\prime}$ evaluated at all points on this cycle (except for the one to the left of $w^{*}$) is strictly smaller than that at $w^{*}$ (Assumption 1). This is sufficient to guarantee that the Jacobian of the $k+1$th iterate GD map has a magnitude strictly less than one; if this is not the case, we can extend the trajectory backwards from $w_{0}$ to $w_{-1}$, $w_{-2}$, …, until satisfied. Local stability implies that if we initialize sufficiently close to the cycle (rather than the minimizer), GD would converge to this cycle instead. ∎
As verifying the assumptions of the implicit function theorem can be technical and tedious, we defer the full proof to Section A.2 and the supplemental materials. In Figure 5, we illustrate that one can easily construct a dataset for different values of $\gamma\in(1,2]$ such that GD converges to cycles. A pattern to note is that the smaller the $\gamma$, the longer period the cycle tends to have, as it takes more steps to move along the “flatter” side of the loss before being kicked back. Note that convergence to $w^{*}$ and stable cycles are not necessarily the only two events that can happen below the $2/\lambda$ step size. If for some dataset there exists a period-$3$ cycle under the step size $\eta=\gamma/\lambda$ for some $\gamma<2$, then a chaotic GD trajectory can also occur for that dataset under the same $\gamma$ (Li and Yorke, 1975).
Figure 5: For each $\gamma$, we construct a one-dimensional dataset such that GD on this problem with the logistic loss converges to a stable cycle under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$. One exception is the last column which seems to suggest GD can even be chaotic when $\gamma<2$. Figures in the first row show the loss evaluated at successive iterates, while the second row shows the power spectral density of the losses at the last $1024$ iterations. Details on each dataset are in Section E.3.
@@ 6 Higher dimensions
In the one-dimensional case, we have shown that under the step size $\eta=\gamma/{\mathcal{L}}^{\prime\prime}(w^{*})$, convergence is guaranteed for all $\gamma\leq 1$, while increasing $\gamma$ beyond $1$ can lead to cycles. In higher dimensions, we no longer have global convergence even when $\gamma$ is small. As we now show, for all $\gamma\in(0,1]$, a cycle can be constructed in $d=2$ using similar techniques. Since we can embed the constructed dataset in higher dimensions, the result trivially extends to any non-linearly-separable classification problem satisfying Assumption 1 of any dimension $d>1$.
@@
For any loss function $\ell$ for which Assumption 1 holds and any $\gamma\in(0,1]$, there exists a 2-dimensional non-separable classification problem of the form 7, on which there exists a GD trajectory under the step size $\eta=\gamma/\lambda$ that converges to a cycle of period $k>1$.
@@
Fix a $\gamma\in(0,1]$, we construct a base dataset corresponding to a loss ${\mathcal{L}}$ such that the minimizer is strictly positive. This base dataset is essentially two 1-dimensional datasets lying independently in the two dimensions. We show that there exists a trajectory that moves almost in a straight line towards the minimizer, as in Figure 6 (left). Let this trajectory be $w_{0},w_{1},\dots$, up to $w_{k-2}$, and $T(w)=w-\eta\nabla{\mathcal{L}}(w)$ be the GD update on ${\mathcal{L}}$.
@@
for some constants $c_{1},c_{2}>0$ and vectors $p$ and $q$ that we will choose later. Specifically, $p$ and $q$ are positioned such that the minimizer of ${\mathcal{L}}_{\text{perturbed}}$ is identical to that of ${\mathcal{L}}$. Let $\left\{\tilde{w}_{t}\right\}_{t}$ be the trajectory of running GD on ${\mathcal{L}}_{\text{perturbed}}$ with the step size $\eta=\gamma/\lambda_{\max}(\nabla^{2}{\mathcal{L}}_{\text{perturbed}}(w^{*}))=\eta$, and the same initialization $\tilde{w}_{0}=w_{0}$.
Observe that the two ReLUs are only activated for all points above the line $w^{\mathsf{\scriptscriptstyle T}}p=0$ and below the line $w^{\mathsf{\scriptscriptstyle T}}q=0$. We design our original trajectory such that neither gets activated until $w_{k-2}$, thus the two trajectories coincide for all $w_{t}$ from $t=0$ to $t=k-2.$ The point $\tilde{w}_{k-2}=w_{k-2}$ crosses over the line defined by $w^{\mathsf{\scriptscriptstyle T}}p=0$, thus activate the first ReLU, kicking the perturbed trajectory in the direction orthogonal to $\tilde{w}_{k-2}$ (as in the middle plot of Figure 6). At this new point $\tilde{w}_{k-1}$, we carefully set $c_{1}$, $c_{2}$, $p$ and $q$ such that the first ReLU becomes inactive while the second activates, and that another GD step takes us back to $w_{0}$.
Figure 6: Illustration of our cycle construction in two dimensions using the logistic loss. The objective (in log scale) in the middle and right contours are created by adding large-magnitude examples to the dataset underlying ${\mathcal{L}}$, depicted on the left. Red stars mark the minimizer $w^{*}$ of the respective losses. Starting at the same $w_{0}$ using the same $\gamma=0.4$, GD on ${\mathcal{L}}$ and ${\mathcal{L}}$ with one set of additional data points both converge to the solution. When we add in the second set of data points, GD converges to a period-$13$ cycle. The trajectory in white marks the first $500$ steps, while the final $20$ steps are marked in red. Details on the synthetic dataset are in Section E.3. A discussion on the basin of attraction of such limit cycle can be found in E.1.
As in the one-dimensional case (Theorem 2), we then use the fact that ${\mathcal{L}}_{\text{perturbed}}$ is the limit of a continuous perturbation for which we can find a valid classification dataset. Stability of the resulting cycle is also straightforward. ∎
While the proof technique is similar to what we used in the one-dimensional case, the two-dimensional case is actually a lot more interesting. If $\gamma$ were to remain in the $(1,2]$ range, then we could have trivially stacked two (or more) one-dimensional constructions to obtain a $d>1$ counterexample. But because we are pushing $\gamma$ below $1$, the unperturbed trajectory on the base dataset cannot cross over 00 in either dimensions as it did in the one-dimensional construction. The difficulty lies in positioning the two kicks albeit the existence of an invariant set in the base dataset.
@@
For any $\ell$ for which Assumption 1 holds, any $\gamma\in(0,2]$ and any $d>2$, there exists a $d$-dimensional non-separable classification problem of the form 7, on which there exists a GD trajectory under the step size $\eta=\gamma/\lambda$ that converges to a cycle of period $k>1$.
@@
The easiest way to construct a counterexample in $d>2$ is to stack multiple lower-dimensional counterexamples. Let such a dataset be $(X^{\prime},y^{\prime})$ with $X^{\prime}$ an $n^{\prime}\times d^{\prime}$ matrix and $d^{\prime}=1$ or $2$. Let its solution be $\bar{w}^{\prime}$, and the corresponding period-$k$ orbit under some $\gamma<2$ be denoted as $\bar{w}_{1}$, $\bar{w}_{2}$, …, $\bar{w}_{k}$. Now consider a new classification problem with features $X=I_{k}\otimes X^{\prime}$, where $I_{k}$ is the $k\times k$ identity matrix, and $\otimes$ is the Kronecker product. The corresponding label vector $y$ is just $y^{\prime}$ repeated $k$ times.
Essentially, we have stacked $k$ independent problems of the original dataset, one across every $d^{\prime}$ dimension(s). As a result, the solution $\bar{w}$ is simply $\bar{w}^{\prime}$ repeated $k$ times. Note that both the gradient an the Hessian are identical to that of the original problem in each dimension, albeit scaled down by a factor of $k$ due to the increase in the number of examples. The Hessian is also (block) diagonal everywhere, so the new step size $\gamma/\lambda$ is the same as the original but scaled up by $k$. This allows the extra $k$ to be cancelled out in the GD update.
Now we can run GD with that step size initialized at $w_{0}\in\mathbb{R}^{kd^{\prime}}$, and as long as $w_{0}$ contains an entry (or two consecutive entries starting at an even index) close enough to a point on the $k$-periodic orbit, the trajectory will converge to a cycle.
In fact, one can mix and match any combination of low-dimensional cycles like so. Choosing the appropriate $\gamma$ will guarantee cycling in at least one of the dimensions. One can even embed a low-dimensional counterexample in an arbitrary problem (of smaller $\lambda$) in the same manner, potentially involving real data. ∎
An interesting observation on the example in the previous proof is that it can actually violate the EoS phenomenon (Cohen et al., 2021). For simplicity, let $d^{\prime}=1$. Suppose we pick a specific initialization $w_{0}$ in the neighborhood of $(\bar{w}_{1},\bar{w}_{2},\dots,\bar{w}_{k})$. Then in each dimension, GD will converge to the original period-$k$ cycle, but off-sync by precisely one time step. That is, each point on this $d=k$-dimensional cycle will be a cyclic permutation of the previous point. Observe that in the original one-dimensional cycle, there is a unique point that attains the highest second derivative, one that is potentially greater than $\lambda=\gamma/\eta$. Combining the fact that the Hessian is diagonal and the $d$-dimensional cyclic orbit is always a permutation, the Hessian will always have its maximum eigenvalue exactly equal to that of the peak, and this peak can be greater than $2/\eta$. As illustrated in Figure 10, we now have an example where the step size used is below the stability threshold, the sharpness converges to a constant (albeit strictly above $2/\eta$), and the loss does not continue to decrease. This seems to contradict the result of Damian et al. (2023), which predicts that GD often self-stabilizes once the sharpness increases above the $2/\eta$. However, their prediction relies on the existence of progressive sharpening, the lack of which in our example might be the reason behind this non-self-stabilizing behaviour.
@@ Appendix A Proofs in one dimension
For all proofs in this section, we assume without loss of generality that $y_{i}=1$ for all $i$. If $y_{i}=-1$, we can simply flip the corresponding $x_{i}$ to $-x_{i}$ to obtain the same objective. We use $T$ to denote the GD map with a constant step size $\eta$
@@ A.1 Convergence under the stable step size
We assume without loss of generality that $w^{*}>0$. If $w^{*}=0$, then fact that ${\mathcal{L}}^{\prime\prime}$ is maximized at 00 implies $T$ is Lipschitz with a constant strictly less than $1$ (because ′′${\mathcal{L}}^{\prime\prime}(w^{*})>0$ when the data is non-separable). Therefore $T$ is a contraction which implies global convergence. If $w^{*}<0$, we can flip the signs of all the $x_{i}$’s and get $w^{*}>0$ as the problem is symmetric about 00. The proof is split into three cases based on $w_{0}$:
@@
$=\underbrace{w^{*}+\eta C<\infty}_{\bar{w}}$ (17)
for some finite $C>0$ due to boundedness of $\ell^{\prime}$. As $\bar{w}$ does not depend on $t$, we can never move arbitrarily far to the right. It remains to establish the iteration complexity.
Case 2: $w_{0}<0$. In this case, the maximum number of iterations to arrive at $w_{\tau}\geq 0$ is
@@
which holds for all $\eta>0$, as ${\mathcal{L}}^{\prime}<0$ but decreasing in magnitude as we increase from $w_{0}$ to 00.
@@
So in this interval, we cross over into ${\mathcal{I}}$ in one step. Combining 16, 18 and A.1, we have that after at most $1+\max\left\{0,\,-\lceil w_{0}\cdot{\mathcal{L}}^{\prime\prime}(w^{*})/{\mathcal{L}}^{\prime\prime}(0)\rceil\right\}$ number of iterations, we will enter ${\mathcal{I}}$, from which convergence is guaranteed at a linear rate. The rate depends on $\eta=1/{\mathcal{L}}^{\prime\prime}(w^{*})$ as well as $\bar{w}$ if we did not initialize in ${\mathcal{I}}$. If $w^{*}<0$, then the initial number of iterations kick in when $w_{0}>0$ and the entire argument holds with a sign flip. This completes the proof. ∎
@@ A.2 Cycle construction below the critical step size
$w_{k-1}<0<w_{k-2}<\dots<w_{1}<w_{0}$ (19)
is possible for some $k>1$, under the step size $\eta$ where $w_{t+1}=T(w_{t})$ for $t=0,1,\dots,k-2$. We also assume that $w_{0}>T(w_{k-1})$. This is valid since if it’s not the case, we can use Lemma 2 to extend the trajectory further to the right as $w_{0}<w_{-1}<\dots<w_{-t}$ until $w_{-t}>T(w_{k-1})$ is satisfied, then relabel the iterations.
@@
$\ell_{\epsilon}(w)\coloneqq\begin{cases}\epsilon^{2}\cdot\ell\left(-\frac{w}{\epsilon^{2}}\right)&\text{if $\epsilon\neq 0$}\\ \max\{-w,\,0\}&\text{if $\epsilon=0$}.\end{cases}$ (20)
We also define the constant
$c\coloneqq\frac{w_{0}-T(w_{k-1})}{\eta}>0,$ (21)
@@
be the corresponding GD map. Starting at $\tilde{w}_{0}=w_{0}$ in 19, let $\tilde{w}_{t+1}=T_{\epsilon}(\tilde{w}_{t})$ for $t=0,1\dots,k$. Since $w^{*}_{\epsilon}$, ${\mathcal{L}}^{\prime}_{\epsilon}$, and ${\mathcal{L}}^{\prime\prime}_{\epsilon}$ are all continuous in $\epsilon$ (by Lemmas 3 and 4), so is $\eta_{\epsilon}$ and $T_{\epsilon}$, with respective limits
@@
as $\epsilon\to 0$. Furthermore, for all $w>0$, $\lim_{\epsilon\to 0}\ell^{\prime}_{\epsilon}(w)=0$. Combined with the fact that original trajectory 19 is positive up to $w_{k-2}$, we also have $\tilde{w}_{t}\to w_{t}$ as $\epsilon\to 0$ for all $t=0,1,\dots,k-1$. We will show that the next step takes us back to $w_{0}$, resulting in a cycle.
@@
$G(\epsilon,w)\coloneqq T^{k}_{\epsilon}(w)-w=\underbrace{T_{\epsilon}\circ\ldots\circ T_{\epsilon}}_{k\text{ times}}(w)-w.$ (24)
@@
$G(0,w_{0})$ $=T_{0}(T^{k-1}_{0}(w_{0}))-w_{0}$
$=T_{0}\left(w_{k-1}\right)-w_{0}$
$=w_{k-1}-\eta L^{\prime}(w_{k-1})+\eta\cdot c-w_{0}$ ($w_{k-1}<0$)
$=T(w_{k-1})+w_{0}-T(w_{k-1})-w_{0}$ (Definition of $c$ in 21)
@@
$\frac{\partial G}{\partial w}(0,w_{0})$ $=\prod_{t=0}^{k-1}\left(1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(T_{0}(\tilde{w}_{t}))\right)-1.$
Note that for all $t=0,1,\dots,k-2$, $\tilde{w}_{t}=w_{t}>w^{*}$ when $\epsilon=0$ and that $w^{*}=w^{*}_{0}$,
@@
$\left|\prod_{t=0}^{k-2}\left(1-\eta_{0}{\mathcal{L}}_{0}^{\prime\prime}(T_{0}(\tilde{w}_{t}))\right)\right|<\left|(1-\gamma)^{k}\right|\leq 1$
@@
$\mu(\epsilon,w)$ $\coloneqq\left|(T_{\epsilon}^{k})^{\prime}(w)\right|=\left|\prod_{t=0}^{k-1}(1-\eta_{\epsilon}{\mathcal{L}}_{\epsilon}^{\prime\prime}(T^{t}_{\epsilon}(w)))\right|,$ (25)
@@
We know that $\mu(0,w_{0})<1$ from earlier, and by continuity there must exist a neighborhood $U$ around $\epsilon=0$ such that $\mu(\epsilon,\omega(\epsilon))<1$ for all $\epsilon\in U$, as $\omega$ is continuous in $\epsilon$. Since $U$ and $I$ are both open, there must exist a nonzero $\epsilon_{0}\in I\cap U$ such that
@@
implying that the trajectory on the perturbed objective with $\epsilon=\epsilon_{0}$ is a stable cycle. Stability ensures that if initialize close enough to $\omega(\epsilon_{0})$ we will converge to this cycle, as it is a stable fixed point of the $T^{k}$ map. A final note is that we can always pick $\epsilon_{0}$ such that $c\cdot\epsilon_{0}^{2}$ is rational, so that the perturbed objective 22 indeed corresponds to a valid classification problem. This completes the proof. ∎
Lemma 1 (Crossing over 00).